2. A bond has a $ 1,900 par value, 10 years to maturity, and 7% annual coupon and sells for $1,800 What is its YTM? Assume YTM remains constant for the next 3 years. What will be its price 3 years from now? a) 7.78 & $5798.23 respectively b) 4.34 & $1826.15 respectively c) 7.78 & $ 1949.17 respectively d) 7.78 & $1822.36 respectively 3. The changes in interest rate can be helpful or hurtful to bondholders. When there is an increase in interest rate, the bond holders are hurt; How are the bondholders hurt when the interest rate decreases. Please explain how this happens (bondholders being hurt when the interest rate decrease).
Answer #1

2: Option D

Using financial calculator

Input: FV = 1900

N =10

PMT = 7%*1900= 133

PV = -1800

Solve for I/Y as 7.78

Hence YTM = 7.78%

Price after 3 years

Using financial calculator

Input: FV = 1900

N =10-3= 7

PMT = 7%*1900= 133

I/Y = 7.78

Solve for PV as -1822.60

Hence price = $1822.60

3: when there is an increase in the interest rates the price of the bond declines. This is because due to an increase in the interest rates the present bonds become less attractive in the market. This reduces the demand for the bonds. Considering the forces of demand and supply due to a drop in the demand the prices of the bonds decline and this reduces the capital gain of the bondholders.